# The information mismatch, and how to fix it

### v1

@@ 1 Introduction
The scenario we have just described is standard in the prevailing medical diagnostic reasoning framework. However, the post-test probability has an error, because sensitivity and specificity do not depend on patient information, or covariates, such as age, symptoms, signs, etc. For example, expositions of the prevailing medical diagnostic reasoning framework, which define sensitivity and specificity in such a way, see, e.g., Ledley and Lusted (1959); Parikh et al. (2008); Trevethan (2017). Although sensitivity and specificity being unadjusted for patient covariates is often taken as a definition in the literature, it is actually an assumption. In particular, it is the assumption that test results are independent of all patient characteristics other than disease status (for a discussion, see Section 1.1.3 of Pearl (2009)).
That such an assumption is questionable has been noted by multiple authors. For example, Dawid (1976) discusses this in light of selection bias, Hlatky et al. (1984) for exercise electrocardiography, and Moons et al. (1997) for coronary disease tests as a function of, e.g., patient covariates obtained from history and physical examination. Guggenmoos-Holzmann and van Houwelingen (2000) criticizes the classical view that sensitivity and specificity are considered “universal constants”(i.e., constant with respect to patient covariates). The diagnosis chapter in Harrell et al. (2001), along with Moons and Harrell (2003), aligns with this and suggests that the role of sensitivity and specificity should be de-emphasized in diagnostic studies. Instead, direct modeling of the post-test probability, adjusted for covariate information, should be performed. This direct modeling is undertaken, for example, by Janssens et al. (2005). The argument for sensitivity and specificity depending on patient covariates is best framed in the language of probability. Therefore, some authors Harrell (2017); Chowdhry (2023) have recommended that test characteristics should be described using conditional probability notation rather than plug-in equations.
@@ 2 Diagnostic problem
Note that the expression in (2) only conditions on the test result, as do most examples from the medical diagnosis literature (see the commonly cited overviews of post-test probability, such as Parikh et al. (2008); Trevethan (2017)). Conditioning only on the test result does not align with clinical practice. The clinician wishes to provide the most patient-specific post-test probability, and this probability is a function not only of the test result but also of the other patient covariates (i.e., other information we have about the patient).
@@ 3 The information mismatch, an example
However, it is rarely the case that $Test\perp X.$ It is unlikely that US is independent of the severity of the DVT, where the severity is taken into account in the Wells’ score based on the tenderness of the leg and the extent of the swelling, since more severe disease is often easier to detect (Harrell et al., 2001). That this independence between the test and covariates does not hold for other medical conditions has been shown in many studies; see, for example, Moons et al. (1997); Hlatky et al. (1984); Prince-Guerra et al. (2021) or the other examples listed in Moons and Harrell (2003); Harrell et al. (2001), and Guggenmoos-Holzmann and van Houwelingen (2000).
@@ 5 Fixing the information mismatch
To truly fix the information mismatch, probabilistic concepts must be described in the language of probability. As discussed in Harrell (2017); Chowdhry (2023), it is crucial that there be less emphasis on terminology, such as ”sensitivity” and its formula in (3) — or, for this matter, “$\text{sensitivity}_{x},$” which could be, against our recommendations, represented as a new formula — and more emphasis on the probability behind these concepts, and how to derive and evaluate probabilistic statements. Ultimately, this is most likely to fix the information mismatch, and will greatly enhance the application of evidence to clinical problems.

### v2

@@ 1 Introduction
The scenario we have just described is standard in the prevailing medical diagnostic reasoning framework. However, the post-test probability has an error, because sensitivity and specificity do not depend on patient information, or covariates, such as age, symptoms, signs, etc. For example, expositions of the prevailing medical diagnostic reasoning framework, which define sensitivity and specificity in such a way, see, e.g., Ledley and Lusted (1959),Parikh et al. (2008), or Trevethan (2017). Although sensitivity and specificity being unadjusted for patient covariates is often taken as a definition in the literature, it is actually an assumption. In particular, it is the assumption that test results are independent of all patient characteristics other than disease status (for a discussion, see Section 1.1.3 of Pearl (2009)).
That such an assumption is questionable has been noted by multiple authors. For example, Dawid (1976) discusses this in light of selection bias, Hlatky et al. (1984) for exercise electrocardiography, and Moons et al. (1997) for coronary disease tests as a function of, e.g., patient covariates obtained from history and physical examination. Guggenmoos-Holzmann and van Houwelingen (2000) criticizes the classical view that sensitivity and specificity are considered “universal constants”(i.e., constant with respect to patient covariates). The diagnosis chapter in Harrell et al. (2001), along with Moons and Harrell (2003), aligns with this and suggests that the role of sensitivity and specificity should be de-emphasized in diagnostic studies. Instead, direct modeling of the post-test probability, adjusted for covariate information, should be performed. This direct modeling is undertaken, for example, by Janssens et al. (2005). The argument for sensitivity and specificity depending on patient covariates is best framed in the language of probability. Therefore, some authors (Harrell, 2017; Chowdhry, 2023) have recommended that test characteristics should be described using conditional probability notation, as in Pagano et al. (2022), rather than plug-in equations.
@@ 2 Diagnostic problem
Note that the expression in (2) only conditions on the test result, as do most examples from the medical diagnosis literature (see the commonly cited overviews of post-test probability, such as Parikh et al. (2008) or Trevethan (2017)). Conditioning only on the test result does not align with clinical practice. The clinician wishes to provide the most patient-specific post-test probability, and this probability is a function not only of the test result but also of the other patient covariates (i.e., other information we have about the patient).
@@ 3 The information mismatch, an example
However, it is rarely the case that $Test\perp X.$ It is unlikely that US is independent of the severity of the DVT, where the severity is taken into account in the Wells’ score based on the tenderness of the leg and the extent of the swelling, since more severe disease is often easier to detect (Harrell et al., 2001). That this independence between the test and covariates does not hold for other medical conditions has been shown in many studies; see, for example, Moons et al. (1997),Hlatky et al. (1984),Prince-Guerra et al. (2021), or the other examples listed in Moons and Harrell (2003),Harrell et al. (2001), and Guggenmoos-Holzmann and van Houwelingen (2000).
@@ 5 Fixing the information mismatch
To truly fix the information mismatch, probabilistic concepts must be described in the language of probability. As discussed in Harrell (2017) and Chowdhry (2023), it is crucial that there be less emphasis on terminology, such as ”sensitivity” and its formula in (3) — or, for this matter, “$\text{sensitivity}_{x},$” which could be, against our recommendations, represented as a new formula — and more emphasis on the probability behind these concepts, and how to derive and evaluate probabilistic statements. Ultimately, this is most likely to fix the information mismatch, and will greatly enhance the application of evidence to clinical problems.